OpenStudy (anonymous):

what value completes the square for the expression x^2+20x

OpenStudy (mertsj):

What is the coefficient of x?

OpenStudy (anonymous):

1?

OpenStudy (mertsj):

The coefficient of x is the number in front of the x

OpenStudy (anonymous):

20

OpenStudy (mertsj):

Divide that number by 2

OpenStudy (anonymous):

10

OpenStudy (mertsj):

Square that number

OpenStudy (anonymous):

100

OpenStudy (mertsj):

That is the number that completes the square.

OpenStudy (anonymous):

ok but whats the reasoning behind it

OpenStudy (mertsj):

The reasoning is that all trinomials that are perfect square trinomials share this property: one-half the coefficient of x, when squared, is the constant term.

OpenStudy (mertsj):

Note these examples:

OpenStudy (mertsj):

\[x^2+6x+9=(x+3)^2\]

OpenStudy (mertsj):

Notice that the coefficient of x is 6. 1/2 of 6 is 3. 3^2 is 9

OpenStudy (anonymous):

oh ok i understand now mind helping me out on a different problem?

OpenStudy (mertsj):

\[x^2-2x+1=(x-1)^2\] Notice that the coefficient of x is -2. 1/2 of -2 is -1 . (-1)^2=1

OpenStudy (mertsj):

What is it?

OpenStudy (anonymous):

Simplify the number using the imaginary unit i: \[\sqrt{-496}\]

OpenStudy (mertsj):

Do you know the definition of i?

OpenStudy (anonymous):

do not but i know its used to remove the negative

OpenStudy (mertsj):

\[i=\sqrt{-1}\]

OpenStudy (mertsj):

\[\sqrt{-496}=\sqrt{16\times 31\times (-1)}=\sqrt{16}\sqrt{31}\sqrt{-1}=4\sqrt{31} i\]

OpenStudy (anonymous):

ok i understand its just difficult for me to gett the two numbers that times to 496 etc

OpenStudy (mertsj):

You could always write the prime factors.

OpenStudy (anonymous):

what about square root of -240 and -169

OpenStudy (mertsj):

240=4 x 60=4 x 4 x 15=16 x 15 169=13^2

OpenStudy (anonymous):

alright thanks last one please

OpenStudy (mertsj):

What is it?

OpenStudy (anonymous):

24x2+3x

OpenStudy (anonymous):

i dont know which 2 numbers add to 24 and multiply to 3

OpenStudy (mertsj):

What are you supposed to do with that? Factor it?

OpenStudy (anonymous):

yes factored form

OpenStudy (mertsj):

There is a common factor of 3x. Factor it out

OpenStudy (anonymous):

so its 3(8x+1)

OpenStudy (mertsj):

Did you factor out 3x as instructed?

OpenStudy (mertsj):

Is 3(8x+1)=24x^2+3x ?????

OpenStudy (anonymous):

dude am lost

OpenStudy (mertsj):

Why don't you just put 3x(8x+1) ? I said to factor out 3x

OpenStudy (anonymous):

so i did do it right

OpenStudy (mertsj):

If you think that \[3x(8x+1)\] is the same as \[3(8x+1)\] then I suppose you could say you did it right but I would not say that.

OpenStudy (anonymous):

ok i forgot x

OpenStudy (anonymous):

what about simplifying expressions (-2+9i)+(-7+3i) i got 33i-59

OpenStudy (mertsj):

14-6i-63i+27^i^2=14-69i-27=-13-69i

OpenStudy (mertsj):

What test are you taking anyway?

OpenStudy (anonymous):

alg 2 midterm review your helping me out a lot and i forgot 14 i wrote 4 but i understand it

OpenStudy (mertsj):

Good

OpenStudy (anonymous):

what about -2+4i/3-i?

OpenStudy (mertsj):

Multiply numerator and denominator by 3+i

OpenStudy (anonymous):

so 6-i/9-i^2?

OpenStudy (mertsj):

What is i^2?

OpenStudy (mertsj):

The whole point is to get rid of the "i" in the denominator.

OpenStudy (mertsj):

\[\frac{-2+4i}{3-i}\times \frac{3+i}{3+i}=\frac{-6-2i+12i+4i^2}{9-i^2}=\frac{-6+10i+4(-1)}{9-(-1)}=\frac{-10+10i}{10}=\]

OpenStudy (mertsj):

\[-1+i\]

OpenStudy (anonymous):

oh alright thanks man for everything

OpenStudy (mertsj):

You had better learn to multiply binomials